### v1

# Learning Minimal NAP Specifications for Neural Network Verification

Abstract.
Specifications play a crucial role in neural network verification. They define the precise input regions we aim to verify, typically represented as L-infinity norm balls. While recent research suggests using neural activation patterns (NAPs) as specifications for verifying unseen test set data, it focuses on computing the most refined NAPs, often limited to very small regions in the input space. In this paper, we study the following problem: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. Finding the minimal NAP specification not only expands verifiable bounds but also provides insights into which neurons contribute to the model’s robustness. To address this problem, we propose several exact and approximate approaches. Our exact approaches leverage the verification tool to find minimal NAP specifications in either a deterministic or statistical manner. Whereas the approximate methods efficiently estimate minimal NAPs using adversarial examples and local gradients, without making calls to the verification tool. This allows us to inspect potential causal links between neurons and the robustness of state-of-the-art neural networks, a task for which existing verification frameworks fail to scale. Our experimental results suggest that minimal NAP specifications require much smaller fractions of neurons compared to the most refined NAP specifications, yet they can significantly expand the verifiable boundaries to several orders of magnitude larger.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
The growing prevalence of deep learning systems in decision-critical applications has elevated safety concerns regarding AI systems, such as their vulnerability to adversarial attacks  (Goodfellow et al., 2015; Dietterich and Horvitz, 2015). Therefore, the verification of AI systems has become increasingly important and attracted much attention from the research community. The field of neural network verification largely follows the paradigm of software verification – using formal methods to verify desirable properties of systems using rigorous mathematical specifications and proofs (Wing, 1990). Most existing works (Katz et al., 2017, 2019; Huang et al., 2017a, 2020; Wang et al., 2021b) utilize $L_{\infty}$ ball specifications for verifying robustness guarantees, i.e., verify if neural networks consistently predict the same label within local neighborhoods ($L_{\infty}$ balls) of reference data points.
Figure 1. The comparison between the NAP specification (depicted by the green region) and the $L_{\infty}$ ball specification (represented by the gray region) in the MNIST dataset. Image $a$ is the reference image, and image $c$ is the closest image to $a$ with an $L_{\infty}$ distance of 0.5294. However, image $c$ cannot be verified using the $L_{\infty}$ ball specification because an adversarial example $b$ can be found at an $L_{\infty}$ distance of 0.2. In contrast, the NAP specification enables verification of inputs that are semantically close despite their geometric distance being far apart, such as image $c$.
While this type of specification provides a fair and effective testbed for evaluating neural networks’ robustness, those verifiable $L_{\infty}$ balls are usually too tight, making them unlikely to cover unseen test set data (Geng et al., 2023). For instance, the $L_{\infty}$ maximum verifiable bounds in the MNIST dataset is smaller than 0.2, whereas the smallest distance among images sharing the same label is greater than 0.5. Thus, narrowing the gap between verifiable regions and real data distribution to further enable verification in real-world settings poses a challenge.
To this end,  Geng et al. (2023) proposes using neural activation patterns (NAPs) as specifications. Neural activation patterns can be viewed as an abstraction of each neuron’s state. For ReLU activation, it can be categorized into an activation state and a deactivation state. Since NAPs represent the intrinsic decision-making processes of neural networks, geologically distant data usually share the same NAPs if they belong to the same class. Compared to conventional local neighborhood specifications, NAP specifications cover much more flexible and potentially larger regions, as illustrated in Figure 1. This allows the verification of unseen real data rather than synthesized data from local neighbourhoods, paving the way for real-world application verification.
However, it is noteworthy that the current approach to computing NAPs relies on a simple statistical method that assumes each neuron contributes to certifying the robustness behaviors of neural networks. Consequently, the computed NAPs are often the most refined ones. However, this is an overly restrictive assumption, as many studies (Frankle and Carbin, 2019; Lu et al., 2019; Liang et al., 2021; Wang et al., 2021a) have revealed that a non-trivial portion of neurons in neural networks may not play a substantial role. This motivates us to address the following challenge: Given a neural network, find a minimal (coarsest) NAP that is sufficient for formal verification of the network’s robustness. This problem is important for the following reasons: i) Minimal NAP specifications correspond to potentially much larger regions in the input space compared to the finest ones, this increases the potential of verification of unseen data; ii) Minimal NAP specifications provide insight into which neurons are responsible for the model making robust predictions, helping us to reveal the black-box nature of neural networks.
To find the minimal NAP specifications, we introduce two basic algorithms: Refine and Coarsen. These algorithms exhaustively check all possible candidates using off-the-shelf verification tools, such as Marabou (Katz et al., 2019). While these approaches provide correctness guarantees, they are not efficient for verifying large neural networks as calls to the verification tool are usually very expensive. To tackle this inefficiency, we propose statistical versions of Refine and CoarsenSample_Refine and Sample_Coarsen that leverage sampling and statistical learning principles to find the minimal NAP specification in a more efficient manner. However, current verification tools cannot scale up to state-of-the-art neural network models, limiting the ability of verification-dependent approaches to compute the minimal NAP specification.
This motivates us to study estimation methods that do not rely on verification tools. To this end, we find that adversarial examples and local gradients can provide valuable information about mandatory neurons, which are the essential building blocks of minimal NAPs. We then translate these findings into two approximate approaches: Adversarial_Prune and Gradient_Search. They can be used together to estimate mandatory neurons of state-of-art neural networks such as VGG-19 (Simonyan and Zisserman, 2014) and inspect possible links between models’ robustness performance and internal neurons, a task that existing verification frameworks struggle to scale for.
Moreover, previous research has claimed that NAP specifications cover significantly larger regions compared to $L_{\infty}$ ball specifications, yet this assertion lacks empirical evidence. In this paper, we address this gap by introducing a simple approach to estimate the volume of regions corresponding to NAP specifications. Furthermore, this method provides us with a rough approximation of the volumetric change from refined NAP specifications to minimal NAP specifications. Our experimental results demonstrate significant volumetric changes, thereby underscoring the efficacy of minimal NAP specifications in real-world verification scenarios. We summarize our contributions as follows:
  • We propose the problem of learning the minimal NAP specification and illustrate its importance in neural network verification. We present two simple approaches that leverage the verification tool with correctness guarantees.
  • We introduce several key concepts related to the problem of learning minimal NAP specifications, including the abstraction function, NAP specification, and mandatory neurons. We demonstrate that the problem can be easily solved with access to mandatory neurons and propose two efficient estimation approaches to learn them.
  • We propose two statistical learning approaches to the minimal NAP specification problem. They leverage sampling and estimation techniques as heuristics to achieve better efficiency.
  • We present a simple approach for estimating the volume of regions corresponding to NAPs. This helps us understand the volumetric change in verifiable space between the minimal NAP specifications and the refined NAP specifications.
  • Our experiments indicate that the minimal NAP specifications involve much smaller fractions of neurons compared to the most refined NAPs. And they significantly expand the verifiable bound by several orders of magnitude.
2. Background
In this section, we introduce basic knowledge and notations of adversarial attacks and neural network verification, with an emphasis on verification using NAP specifications. This may help readers better understand the importance of learning minimal NAP specifications.
2.1. Neural Networks for Classification Tasks
In this paper, we focus on feed-forward ReLU neural networks. Generally speaking, a feed-forward network $N$ is comprised of $L$ layers, where each layer performs a linear transformation followed by a ReLU activation. We denote the pre-activation value and post-activation value at the $l$-th layer as $z^{(l)}(x)$ and $\hat{z}^{(l)}(x)$, respectively. The $l$-th layer computation is expressed as follows: $z^{(l)}(x)=\mathbf{W}^{(l)}\hat{z}^{(l-1)}(x)+\mathbf{b}^{(l)}$, $\hat{z}^{(l)}(x)=\mathbf{ReLu}(z^{(l)}(x))$, with $\mathbf{W}^{(l)}$ being the weight matrix and $\mathbf{b}^{(l)}$ representing the bias for the $l$-th layer. We denote the number of neurons in the $l$-th layer as $d_{l}$, and the $i$-th neuron in layer $l$ as $N_{i,l}$. The pre-activation value and post-activation value of $N_{i,l}$ at input $x$ is computed by $z_{i}^{(l)}(x)$ and $\hat{z}_{i}^{(l)}(x)$. The network can also be viewed as a function $\mathbf{F}^{<N>}:=\mathbb{R}^{d_{0}}\rightarrow\mathbb{R}^{d_{L}}$ such that $\mathbf{F}^{<N>}(x):=z^{(L)}(x)$, where $\mathbf{F}^{<N>}_{i}(x):=z_{i}^{(L)}(x)$ represents the output of $i$-th neuron in the last layer. We will omit $N$ when the context is clear. In a multi-class classification task with $C$ as the set of classes, when provided with an input $x$, the neural network predicts class/label $c\in C$ if the output of the $c$-th neuron $\textbf{F}_{c}(x)$ is the highest among the neurons in the $L$-th layer.
2.2. Adversarial Attacks and Neural Networks Verification Problem
Given a neural network $\mathbf{F}$ and a reference point $x$, adversarial attacks aim to search for a point $x^{\prime}$ that is geometrically close to the reference point $x$ such that $x^{\prime}$ and $x$ belong to different classes. Here, we use the canonical specification, that is, we want to search in the local neighborhoods ($L_{\infty}$ norm balls) of $x$, formally denoted as $B(x,\epsilon):=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}$, where $\epsilon$ is the radius. For certain $\epsilon$, given we know $x$ belongs to class $j$, we say an adversarial point is found if:
(1) $\exists x^{\prime}\in B(x,\epsilon)\quad\exists i\in C\quad\mathbf{F}_{i}(x^{\prime})-\mathbf{F}_{j}(x)>0$
In practice, the change from the original data $x$ to adversarial data $x^{\prime}$ should be imperceptible, so they are more likely to be recognized as the same class/label from a human perspective. There are also metrics other than the $L_{\infty}$ norm to represent the ”similarity” between $x$ and $x^{\prime}$, such as the $L_{0}$ and $L_{2}$ norms (Xu et al., 2020). However, almost all of them fall into the local neighborhood specification paradigm. This is different from the NAP specification, as we will discuss later.
Neural networks have been shown to be vulnerable to adversarial attacks. Even imperceptible changes could easily alter the predicted outcome. This highlights the importance of neural network verification. The neural network verification problem can be seen as the flip side of adversarial attacks. When it can be formally proven that no adversarial point exists in $B(x,\epsilon)$, the neural network verification problem is considered solved. Formally, we want to verify:
(2) $\forall x^{\prime}\in B(x,\epsilon)\quad\forall i\neq j\quad\mathbf{F}_{j}(x^{\prime})-\mathbf{F}_{i}(x^{\prime})>0$
For a simpler presentation, we assume that $\mathbf{F}(x)$ is a binary classifier. For any given specification, $\mathbf{F}(x)\geq 0$ indicates that the model is verified; otherwise, we can find an adversarial example. Solving such a problem is known to be NP-hard (Katz et al., 2017), and current state-of-the-art approaches rely on specialized solvers. The difficulty escalates significantly as the number of neurons in neural networks increases, limiting successful verifications to only very small neural networks. Therefore, achieving scalability in verification remains an ongoing challenge.
2.3. Neuron Abstractions and Neural Activation Patterns
To better discuss robustness verification using NAP specifications, we first introduce the relevant concepts of neuron abstractions, neuron abstraction functions, and neural activation patterns.
Neuron Abstractions
[…]
Definition 2.1 (Neuron Abstraction Function).
Given a neural network $N$ and the abstraction state set $\mathbb{S}$. A neuron abstraction function is the mapping $\mathcal{A}:N\rightarrow\mathbb{S}$. Formally, for an arbitrary neuron $N_{i,l}$ where $0\leq i\leq d_{l}$ and $11\leq l\leq L-$ 1, the function abstracts $N_{i,l}$ to some state $s_{k}\in\mathbb{S}$, i.e., $\mathcal{A}(N_{i,l})=s_{k}$.
The above characterization of neuron abstraction , it does not instruct us on how to perform binary abstraction in the absence of neuron values. Thus, we include the input(s) $x$ or $X$ as a new parameter in $\mathcal{A}$, and it is omitted when the context is clear. Here, we discuss three types of $\mathcal{A}$: the unary abstraction function function $\dot{\mathcal{A}}$, the binary abstraction function $\ddot{\mathcal{A}}$, and the statistical abstraction function $\widetilde{\mathcal{A}}$.
The unary $\dot{\mathcal{A}}$ function
[…]
(3) $\dot{\mathcal{A}}(N_{i,l},x)=\mathbf{*}$
The binary $\ddot{\mathcal{A}}$ function
When an input $x$ is passed through an internal neuron $N_{i,l}$, we can easily determine the binary abstraction state of $N_{i,l}$ based on its post-activation value $\hat{z}_{i}^{(l)}(x)$. This motivates us to define the binary abstraction function as follows:
(4) $\ddot{\mathcal{A}}(N_{i,l},x)=\begin{cases}\mathbf{0}&\text{if }\hat{z}_{i}^{(l)}(x)=0\\ \mathbf{1}&\text{if }\hat{z}_{i}^{(l)}(x)>0\\ \end{cases}$
The statistical $\widetilde{\mathcal{A}}$ function
The binary $\ddot{\mathcal{A}}$ function is limited in calculating a single input, and meets a challenge when multiple inputs come into play, as two distinct inputs may lead to disagreements in a neuron’s abstraction state. We then approach this problem statistically by introducing the $\widetilde{\mathcal{A}}$ function defined as follows:
(5) $\widetilde{\mathcal{A}}(N_{i,l},X)=\begin{cases}\mathbf{0}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{0},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{1}&\text{if }\frac{|\{x_{j}|\ddot{\mathcal{A}}(N_{i,l},x_{j})=\mathbf{1},\text{ }x_{j}\in X\}|}{|X|}\geq\delta\\ \mathbf{*}&\text{otherwise}\\ \end{cases}$
where $\delta$ is a real number from $[0,1]$, and $X$ represents a set of inputs, i.e., $X:=\{x_{1},x_{2},...,x_{n}\}$. The parameter $\delta$ represents an error ratio to accommodate possible disagreements on neuron states between multiple inputs. Intuitively, the introduction of $\delta$ allows multiple inputs $x_{1},...,x_{n}$ to vote on a neuron’s state. For instance, when $\delta$ is set to 0.99, then 99% or more of the inputs must agree that a neuron is activated for the neuron to be in the $\mathbf{1}$ state.
(a) A simple 2x2 fully connected neuron network.
(b) A subset of the NAP $\mathcal{P}$ family forms a binary tree. Each child subsumes its parent, and the root is the most abstracted NAP. The leaf node represents the most refined NAPs, where each neuron is either $\mathbf{1}$ or $\mathbf{0}$. From the root, the tree is spanned by abstracting each neuron in an order of $N_{0,1},N_{1,1},N_{0,2},N_{1,2}$. Setting a different order will create a different tree of NAPs.
Figure 2. A simple 2x2 fully connected neural network and a subset of its NAP in a binary tree structure.
Definition 2.2 (Neural Activation Pattern).
Given a neural network $N$ and a neuron abstraction function $\mathcal{A}$. A neural activation pattern (NAP) $P$ is a tuple that consists of the abstraction state of all neurons in $N$. Formally, $P:=\langle\mathcal{A}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$, also denoted as $\mathcal{A}(N)$. The neuron $N_{i,l}$’s abstraction state specified by NAP $P$ is represented as $P_{i,l}$, i.e., $P_{i,l}=\mathcal{A}(N_{i,l})$.
Partially ordered NAP
We denote the power set of NAPs in $N$ as $\mathcal{P}$. The number of all possible NAPs in $N$ scales exponentially as the number of neurons increases. For such a large set, if we aim to find the minimal NAP — the central problem in this work, we first have to establish an order so that NAPs can be compared. To this end, we define the following partial order:
Definition 2.3 (Partially ordered NAP).
For any given two NAPs $P,P^{\prime}\in\mathcal{P}$, we say $P^{\prime}$ subsumes $P$ if, for each neuron $N_{i,l}$, its state in $P$ is an abstraction of that in $P^{\prime}$. Formally, this can be defined as:
(6) $P^{\prime}\preccurlyeq P\text{ }\iff\text{ }P^{\prime}_{i,l}\preceq P_{i,l}\text{ }\text{ }\forall N_{i,l}\in N$
Moreover, two NAPs $P,P^{\prime}$ are equivalent if $P\preccurlyeq P^{\prime}$ and $P^{\prime}\preccurlyeq P$.
To give a concrete example, Figure 1(b) depicts a subset of NAPs of a simple neural network consisting of 2 hidden layers and 4 neurons, as presented in Figure 1(a).
[…]
precedes-or-equals
[…]
1011
[…]
. This is because these NAPs sit on different branches formed by splitting certain ReLUs.
Region Outlined by NAPs
One key requirement for verification specification is its ability to represent a specific region within the input space.
[…]
Figure 3. Linear regions are shattered by the simple 2X2 neural network. Each linear region corresponds to a most refined NAP, but not necessarily vice versa. More abstracted NAPs are formed by ignoring lines/hyperplanes created by neurons.
2.4. NAP Specifications for Robustness Verification
This subsection illustrates how NAP specifications can be utilized for robustness verification. We discuss how the abstraction states of neurons can serve as signatures of distinct classes with the introduction of the class NAP definition.
Definition 2.4 (Class NAP).
In a classification task with the class set $C$, for any class $c\in C$, a class NAP $P^{c}$ is a NAP comprising abstract states outputted by an abstraction function given $X_{c}$, where $X_{c}$ denotes the set of inputs belonging to class $c$. Formally, $P^{c}:=\langle\mathcal{A}(N_{i,l},X_{c})\mid N_{i,l}\in N,\mathcal{A}\in\{\widetilde{\mathcal{A}},\dot{\mathcal{A}}\}\rangle$. The power set of $P^{c}$ is denoted as $\mathcal{P}^{c}$.
Recall that robustness verification can be thought of as proving that no adversarial examples exist in the local neighbourhoods of some reference point $x$. Intuitively, for NAP specifications, it is equivalent to show that no adversarial examples exist in the class NAP $P^{c}$; in other words, inputs exhibiting the NAP $P^{c}$ must be predicted as class $c$. Geng et al. (2023) argues that class NAPs must satisfy several essential requirements to qualify as NAP specifications. And finding qualified NAP specifications implies the underlying robustness problem is verified. We formally frame these requirements into the following properties:
The non-ambiguity property
Since we want class NAPs to serve as certificates for a certain class, they must be distinct from each other. Otherwise, there could exist an input that exhibits two class NAPs, which leads to conflicting predictions. Formally, we aim to verify the following:
[…]
Pc2
[…]
From a geological perspective, there must be no overlaps between class NAPs. In other words, it is also equivalent to verifying:
$\forall c_{1},c_{2}\in C\text{ s.t. }c_{1}\neq c_{2}\quad R_{P^{c_{1}}}\bigcap R_{P^{c_{2}}}=\emptyset$
We empirically observe that successful models, particularly neural networks that have achieved high accuracy in classification tasks, naturally exhibit this property. Because successful models tend to avoid confusion in predictions, their class NAPs are usually very distinct from each other.
The NAP robustness property
To serve as NAP specifications, class NAPs $P^{c}$ ensure that if an input exhibits it, i.e., $\mathcal{A}(N,x)\preccurlyeq P,\mathcal{A}\in\{\ddot{\mathcal{A}},\dot{\mathcal{A}}\}$, the input must be predicted as the corresponding class $c$. Formally, we have:
(10) $\forall x\in R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
in which
(11) $R_{P^{c}}=\{x\mid{\mathcal{A}(N,x)\preccurlyeq P^{c}}\}$
In contrast to canonical $L_{\infty}$ norm balls, class NAPs are more flexible in terms of size and shape. Additionally, there is no need to specify a reference point, since the locations of potential reference points are also encoded by class NAPs. However, it is possible that no class NAP $P^{c}$ in $\mathcal{P}^{c}$ satisfies this property. This can be mitigated by meeting the subsequent weaker property.
The NAP-augmented robustness property
Instead of relying solely on class NAPs as specifications, local neighbours can still be employed in conjunction for verification. This hybrid form of specification has several advantages: 1) It narrows down the scope of verifiable regions when no class NAPs can meet the NAP robustness property; 2) NAP constraints essentially fix ReLU states, refining the search space for verification tools; 3) It focuses on the verification of valid test inputs rather than adversarial examples. Formally, this property can be stated as:
(12) $\forall x^{\prime}\in B(x,\epsilon)\bigcap R_{P^{c}}\quad\forall k\in C\text{ s.t. }k\neq c\quad\mathbf{F}_{c}(x)-\mathbf{F}_{k}(x)>0$
in which
(13) $B(x,\epsilon)=\{x^{\prime}\mid||x-x^{\prime}||_{\infty}\leq\epsilon\}\quad R_{P^{c}}=\{x^{\prime}\mid{\mathcal{A}(N,x^{\prime})\preccurlyeq P^{c}}\}$
To summarize, we state that a class NAP can serve as a NAP specification if it satisfies either the NAP robustness property or the NAP-augmented robustness property. Clearly, the former property is stronger, and it is possible that we can’t find a class NAP $P^{c}$ in $\mathcal{P}^{c}$ to satisfy this property. Fortunately, we can always find NAPs that satisfy the latter property by narrowing the verifiable region using additional $L_{\infty}$ norm ball specifications. Moreover, NAP specifications can be easily encoded with most of the state-of-the-art verification tools such as Marabou. In this paper, we will not dive into NAP encodings. Interested readers can refer to Geng et al. (2023) for more details.
3. The minimal NAP specification problem
In this section, we formulate the problem of learning minimal NAP specifications and present two naive approaches for solving this problem. Since our approaches require interactions with verification tools, we begin by introducing relevant notations describing the relationships between NAPs and verification tools.
Figure 4. A DAG representing all possible NAPs, with green nodes signifying they can serve as a specification, while red nodes indicate they cannot. Within each node, a black circle means the corresponding neuron is abstracted by $\widetilde{\mathcal{A}}$, whereas the white circle means it is abstracted by $\dot{\mathcal{A}}$. Thus, the top node represents the coarsest NAP, while the bottom node entails the most refined one. Our goal is to find the NAP that resides in the green nodes while its size is minimal.
3.1. Problem Formulation
Let $(\mathcal{P}^{c},\preccurlyeq)$ be a partially ordered set corresponding to a family of class NAPs regarding some class $c\in C$. For simplicity, we omit the superscript $c$ and refer to class NAPs simply as NAPs when the context is clear. We assume we have access to a verification tool, $\mathcal{V}:\mathcal{P}\to\{0,1\}$, which maps a class NAP $P\in\mathcal{P}$ to a binary set. Here, $\mathcal{V}(P)=1$ denotes a successful verification of the underlying robustness query, while 0 indicates the presence of an adversarial example. From an alternative perspective, $\mathcal{V}(P)=1$ also signifies that $P$ is a NAP specification, i.e., it satisfies NAP(-augmented) robustness properties; whereas $\mathcal{V}(P)=0$ implies the opposite.
It is not hard to see that $\mathcal{V}$ is monotone with respect to the NAP family $(\mathcal{P},\preccurlyeq)$. Given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P^{\prime})=1$, it follows that $\mathcal{V}(P)=1$. However, given $P\preccurlyeq P^{\prime}$ and $\mathcal{V}(P)=1$, we cannot determine $\mathcal{V}(P^{\prime})$. In other words, refining a NAP (by increasing the number of neurons abstracted to $\mathbf{0}$ or $\mathbf{1}$) can only enhance the likelihood of successful verification of the underlying robustness query.
Definition 3.1 (The minimal NAP specification problem).
Given a family of NAPs $\mathcal{P}$ and a verification tool $\mathcal{V}$, the minimal NAP specification problem is to find a NAP $P$ such that
$\underset{P\in\mathcal{P},\mathcal{V}(P)=1}{\arg\min}|P|$
where $|P|$ is defined as the size of $P$, representing the number of neurons that are abstracted to the $\mathbf{0}$ or $\mathbf{1}$, i.e., |$|\{N_{i,l}\mid P_{i,l}=\mathbf{0}\text{ or }\mathbf{1}\}|$. The (largest) minimal NAP specification’s size is denoted as $s$.
When $P$ is minimal, it implies that for any NAP $P^{\prime}$ that is strictly more coarse than $P$, $\mathcal{V}(P^{\prime})=0$. Formally, this can be expressed as: $\forall P\preccurlyeq P^{\prime},P^{\prime}\neq P:\mathcal{V}(P^{\prime})=0$. Thus, there could exist multiple minimal NAP specifications. In such cases, we only need to choose one of them. On the other hand, it is possible that even the most refined NAPs cannot verify the robustness query. In such cases, we claim that no minimal NAP specifications exist. Additionally, since the computation using verification tools is usually expensive, we are interested in methods that efficiently find a minimal NAP specification, i.e., minimizing the number of calls to $\mathcal{V}$.
We present two naive approaches to solving the problem. Since these approaches involve refining and coarsening the NAP, we will first formally define these two actions.
Refine and Coarsen Action of NAP
Recall our definition of NAP, the coarsest NAP is the one using $\dot{\mathcal{A}}$ to abstract each neuron in $N$. We denote this NAP as $\dot{P}:=\langle\dot{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\dot{P}$ is the smallest NAP with $|\dot{P}|=0$. In addition, we define the most refined NAP as the one that applies $\widetilde{\mathcal{A}}$ to abstract each neuron in $N$. We denote it as $\widetilde{P}:=\langle\widetilde{\mathcal{A}}(N_{i,l})\text{ }|\text{ }N_{i,l}\in N\rangle$. $\widetilde{P}$ is the largest NAP with size with $|\widetilde{P}|\leq|N|$. Clearly, $\widetilde{P}\preccurlyeq\dot{P}$. Given any NAP $P$, if we want to refine $P$ through a specific neuron $N_{i,l}$, we apply the $\widetilde{\mathcal{A}}$ function to $N_{i,l}$. We denote this refinement action as $\widetilde{\Delta}(N_{i,l})$. This action will either increase or leave unchanged $|P|$. Similarly, we denote the coarsen action as $\dot{\Delta}(N_{i,l})$. This action will either decrease or leave unchanged $|P|$. We present the semantics of these actions as follows:
               $P\vdash\cdot\Downarrow P$
  $P\vdash\Delta_{1}\Downarrow P^{\prime}$         $P^{\prime}\vdash\Delta_{2}\Downarrow P^{\prime\prime}$              $P\vdash\Delta_{1};\Delta_{2}\Downarrow P^{\prime\prime}$
                           $P_{i,l}=\mathbf{*}$ Refine              $P\vdash\widetilde{\Delta}(N_{i,l})\Downarrow\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$                                 $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ Coarsen                $P\vdash\dot{\Delta}(N_{i,l})\Downarrow\dot{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$
where $P,P^{\prime},P^{\prime\prime}$ are from the NAP family $\mathcal{P}$, and $N$ is the underlying neural network. $\Delta_{1}$ and $\Delta_{2}$ are any two sequences of refine and coarsen actions.
3.2. The Refine Approach
Conceptually, the Refine approach iteratively increases the number of refined neurons in NAP $P$ until $\mathcal{V}(P)=1$, i.e., $P$ is able to prove the underlying robustness query. In other words, we gradually increase the size parameter $k$ and iterate over each NAP $P$ of size $k$ to check if $\mathcal{V}(P)=1$, as illustrated in Algorithm 1. To determine if a solution to the problem exists, we first check if the most refined NAP can succeed in verification. We proceed to iterative refinement only if $\mathcal{V}(\widetilde{P})=1$. However, the algorithm is not efficient and requires $12^{|N|}-$ 1 calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.2. Please refer to the proof in Appendix A. Therefore Refine is only practical when the search space of the NAP family $\mathcal{P}$ is small.
Theorem 3.2.
The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
3.3. The Coarsen Approach
In contrast to the Refine approach, the Coarsen approach starts from the most refined NAP and then gradually coarsens it. We first check if the problem is well-defined by verifying if the most refined NAP $\widetilde{P}$ succeeds in verification. Then, for each neuron, we attempt to coarsen it using $\dot{\mathcal{A}}$; if the resulting NAP no longer verifies the query, we refine it back using $\widetilde{\mathcal{A}}$; otherwise, we keep the coarsened NAP. We describe the above procedure in Algorithm 2. The algorithm could require $|N|$ calls to $\mathcal{V}$ in the worst case, as proven in Theorem 3.3. Please refer to the proof in Appendix A.
Theorem 3.3.
The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
Input: The neural network $N$
Output: A minimal NAP specification $P$
1 Function Refine(N)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
3       if $\mathcal{V}(P)==0$ then
             return $None$ ;   /* Return None if even the most refined NAP fails verification */
4
5       else
6             for $k$ from 1 to $|N|$ do
                   $P\leftarrow\dot{\mathcal{A}}(N)$   /* Refinement starts from the coarsest NAP $\dot{P}$/
                   $K\_comb\leftarrow$ $Pick\_K\_Neurons(k,N)$   /* Returns all combinations of size k */
7                   for $comb\textbf{ in }K\_comb$ do
8                         for $N_{i,l}\textbf{ in }comb$ do
                               $P\leftarrow\widetilde{\Delta}(N_{i,l})$  /* Refine neurons in the chosen combination */
9
10                        if $|\mathcal{V}(P)|==1$ then
                               return $P$   /* Found the minimal NAP */
11
12
13
14
15
Algorithm 1 refine
Input: The neural network $N$
Output: A minimal NAP specification $P$
1 Function Coarsen(N)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$
3       if $\mathcal{V}(P)==0$ then
             return $None$ ;   /* Return None if the most refined NAP fails verification */
4
5       else
6             for $N_{i,l}$ in $N$ do
                   $P\leftarrow\dot{\Delta}(N_{i,l})$ ;   /* Try to abstract this neuron using $\dot{\Delta}$/
7                   if $\mathcal{V}(P)==0$ then
                         $P\leftarrow\widetilde{\Delta}(N_{i,l})$ ;   /* Refine the neuron back if the verification fails */
8
9
10            return $P$;
11
Algorithm 2 Coarsen
4. Verification Free Approaches
While the Refine and Coarsen algorithms can find minimal NAP specifications with correctness guarantees, their inefficiency poses challenges with the verification of large neural networks. To address this issue, we introduce two efficient approaches for estimating minimal NAP specifications without requiring expensive calls to the verification tool. Unlike verification-based approaches, these estimation methods are deeply linked to mandatory neurons, a key concept in dissecting the minimal NAP specification problem, as discussed below.
Definition 4.1 (Mandatory neuron).
[…]
$M=\{N_{i,l}\mid P_{i,l}\in\{\mathbf{0},\mathbf{1}\},P\textit{ is minimal}\}$
Note that $M$ is the union of the set of mandatory neurons from all minimal NAP specifications. It follows that $|M|\geq s$, where $s$ denotes the size of the largest minimal NAP specification.
The minimal NAP specification problem can be solved trivially if we gain access to $M$. Thus, our verification-free approaches are designed to determine mandatory neurons and estimate $M$. To better understand these our approaches, we first discuss the properties of mandatory neurons. Recall that verifying a robustness query given a NAP specification $P$ is equivalent to showing that $\mathbf{F}(x)\geq 0$ for input $x$ in region $R_{P}$. Thus, the necessary conditions of a mandatory neuron $N_{i,l}$ can be written as follows:
  1. (1)
    If $N_{i,l}$ is in state $\mathbf{0}$, it implies when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)\geq 0$. In addition, $\exists x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)<0$.
  2. (2)
    If $N_{i,l}$ is in state $\mathbf{1}$, it implies $\forall x\text{ s.t. }\hat{z}_{i}^{(l)}(x)>0,\mathbf{F}(x)\geq 0$. In addition, when $\hat{z}_{i}^{(l)}(x)=0,\mathbf{F}(x)<0$.
[…]
4.1. Adversarial-guided Prune
We first introduce Adversarial_Prune to identify mandatory neurons. Intuitively, it attempts to show a neuron $N_{i,l}$ is mandatory by providing an adversarial example $x^{\prime}$ exists.
When an adversarial example $x^{\prime}$ is found, it immediately indicates that the NAP $\ddot{\mathcal{A}}(N,x^{\prime})$ fails the verification, i.e., $\mathcal{V}(\ddot{\mathcal{A}}(N,x^{\prime}))=0$. Moreover, it also implies that any NAP subsumed by $\ddot{\mathcal{A}}(N,x^{\prime})$ fails verification. For instance, suppose an adversarial example $x^{\prime}$ is found for a simple one-layer four-neuron neural network and $\ddot{\mathcal{A}}(N,x^{\prime})$ is $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{0}\rangle$. We can infer that NAPs like $\langle\mathbf{1},\mathbf{0},\mathbf{1},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{0}\rangle$, and $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$ fail the verification. This information is particularly useful when determining if a neuron is mandatory. For example, if we know that NAP $P:=\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{1}\rangle$ is a specification, i.e., $\mathcal{V}(P)=1$, then we can easily deduce that the fourth neuron $N_{4,1}$ is mandatory. This is because that coarsening the fourth neuron would expand $P$ to $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{*}\rangle$, which would include the adversarial example $x^{\prime}$ and thus fail the verification, as illustrated in Figure 4(a). It is evident that the neuron where $P$ and $\ddot{\mathcal{A}}(N,x^{\prime})$ disagrees must be mandatory.
However, when the two NAPs disagree on multiple neurons, things become a little bit different. Suppose the NAP specification $P$ is $\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle$, i.e., $\mathcal{V}(\langle\mathbf{1},\mathbf{1},\mathbf{*},\mathbf{1}\rangle)=1$. We know $\mathcal{V}(\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle)=0$ by the adversarial example $x^{\prime}$. In this case, if we coarsen the second and fourth neurons, $N_{2,1}$ and $N_{4,1}$, the NAP specification will expand to $\langle\mathbf{1},\mathbf{*},\mathbf{*},\mathbf{*}\rangle$, which will cover the $\langle\mathbf{1},\mathbf{0},\mathbf{*},\mathbf{0}\rangle$, thus failing the verification. In this case, $N_{2,1}$ and $N_{4,1}$ could both be mandatory neurons or either one of them is mandatory, as illustrated in Figures 4(b), 4(c), 4(d). So, we simply let $\{N_{2,1},N_{4,1}\}$ be the upper bound of mandatory neurons (learned from $x^{\prime}$). Formerly, given a NAP $P$, we say a neuron $N_{i,l}$ is in the upper bound of mandatory neurons $M$ if satisfies the following condition:
  1. (1)
    $N_{i,l}$ must be in the binary states, i.e., $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$
  2. (2)
    There exists $x^{\prime}$ such that $\ddot{\mathcal{A}}(N_{i,l},x^{\prime})$ XORs with $P_{i,l}$, i.e., $\exists x^{\prime}\text{ such that }\ddot{\mathcal{A}}(N_{i,l},x^{\prime})\oplus P_{i,l}=1$
(a) $N_{4,1}$ is mandatory.
(b) $N_{4,1},N_{2,1}$ are mandatory.
(c) $N_{4,1}$ is mandatory.
(d) $N_{2,1}$ is mandatory.
Figure 5. Geometric interpretation of NAPs on mandatory neurons. The first subfigure represents the case when the two NAPs disagree on $N_{4,1}$. The other three subfigures represent the three cases when the two NAPs disagree on $N_{4,1},N_{2,1}$. Regions colored green pass verification, whereas red indicates an adversarial example exists. Here, we omit states for $N_{1,1},N_{3,1}$ in those NAPs for simplicity.
The field of adversarial attacks has been extensively studied, offering a wealth of methods that we can leverage. These approaches are usually computationally efficient, making it easy to access a large collection of adversarial examples. Therefore, we can compute the upper bound of mandatory neurons efficiently by simply taking the union of upper bounds learned for each adversarial example, as illustrated in Algorithm 3.
Input: The neural network $N$, the input dataset $X$
Output: A collection of mandatory neurons
1 Function Adversarial_Prune($N$, $X$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $x_{j}$ in $X$ do
3             $x^{\prime}_{j}\leftarrow Adversarial\_Attack(x_{j})$ for $N_{i,l}$ in $N$ do
4                   if $P_{i,l}\in\{\mathbf{0},\mathbf{1}\}$ and $\ddot{\mathcal{A}}(N_{i,l},x^{\prime}_{j})\oplus P_{i,l}$ then
                         $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Create mandatory neurons for $x^{\prime}_{j}$/
5
6
7      return $Mand$
Algorithm 3 Adversarial_Prune
4.2. Gradient-guided Search
We introduce another approach called Gradient_Search to identify mandatory neurons. Similar to Adversarial_Prune, Gradient_Search avoids costly interactions with the verification tool. Instead, it leverages gradient estimations to analyze the structure of $\mathbf{F}$.
Recall our definition of a neural network in Section 2.1, where the output $\mathbf{F}(x)$ can be seen as a function of the post-activation value of any internal neuron $\hat{z}_{i}^{(l)}$. We denote this function as $\mathbf{F}(\hat{z}(x))$, omitting $i$ and $l$ for simplicity. The function $\mathbf{F}(\hat{z})$ is non-linear and operates within an input range of $\hat{z}\in[0,+\infty)$. It is worth noting that the gradient $\frac{\partial\mathbf{F}}{\partial\hat{z}}$ can provide valuable insights into the local structure of $\mathbf{F}$. This insight may help determine the necessity condition of mandatory neurons by checking if there exists a $\hat{z}$ such that $\mathbf{F}(\hat{z})<0$ for the corresponding neuron. To be more specific, given some sampled data $\hat{z}_{1},\ldots,\hat{z}_{n}$, we can compute their corresponding function values $\mathbf{F}(\hat{z}_{1}),\ldots,\mathbf{F}(\hat{z}_{n})$ along with their gradients $\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{1}},\ldots,\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{n}}$. Then if there exists a $\hat{z}_{j}$ such that $|\mathbf{F}(\hat{z}_{j})|$ is sufficiently small, yet the norm of its gradient $|\frac{\partial\mathbf{F}}{\partial\hat{z}}|_{\hat{z}_{j}}|$ is significantly large, then it is very likely that $\mathbf{F}$ will go below zero nearby, as illustrated in Figure 6. Then, if we happen to know that the corresponding neuron $N_{i,l}$ is in state $\mathbf{0}$, it is highly likely that $N_{i,l}$ is mandatory. Conversely, if a neuron $N_{i,l}$ is in state $\mathbf{1}$, and we sample an $x$ such that $\hat{z}_{i}^{(l)}=0$ and $\mathbf{F}(x)<0$, we immediately recognize that the neuron is mandatory. Algorithm 4 presents pseudocode for estimating whether a neuron is mandatory based on the discussions above.
Figure 6. The gradient provides insights into the local structure of $\mathbf{F}$. When the gradient is large and the function value is small, it strongly suggests that $\mathbf{F}$ will go below zero nearby.
Input: The neural network $N$, the input dataset $X$, and thresholds $\beta$, $\gamma$
Output: A collection of mandatory neurons
1 Function Gradient_Search($N$, $X$, $\beta$, $\gamma$)
2       $Mand\leftarrow\emptyset$ ; $P\leftarrow\widetilde{\mathcal{A}}(N)$ for $N_{i,l}$ in $N$ do
3             for $x_{j}$ in $X$ do
4                   $\hat{z}_{j}\leftarrow\hat{z}_{i}^{(l)}(x_{j})$ if $\hat{z}_{j}>0$  then
5                         if  $|\mathbf{F}(\hat{z}_{j})|<\beta$ and $|\frac{\partial\mathbf{F}}{\partial\hat{z}_{i}^{(l)}}|_{\hat{z}_{j}}|>\gamma$ and $P_{i,l}==\mathbf{0}$ then
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{0}$, can’t coarsen */
6
7                   else
8                         if $\mathbf{F}(\hat{z}_{j})<0$ and $P_{i,l}==\mathbf{1}$ then
                               $Mand\leftarrow Mand\cup\{N_{i,l}\}$   /* Neuron is in state $\mathbf{1}$, can’t coarsen */
9
10
11
12
13      return $Mand$
Algorithm 4 Gradient_Search
Both the estimation method Gradient_Search and Adversarial_Prune can be used together. Taking the intersection of their estimated results may provide us with a better overall estimate. In addition, the estimated mandatory neurons can serve as an initial starting point in our statistical versions of the Refine and Coarsen approaches introduced in Section 5.
Our experimental results regarding these two estimation approaches indicate a notable trend: neurons from deeper layers are more likely to be mandatory. This observation aligns with commonly-held beliefs that deeper neurons are generally linked to high-level feature representation, and thus play a more important role than shallow neurons in final classification decisions. This observation leads us to a valuable heuristic: prioritizing deeper neurons over shallower ones as we iterate through neurons in our implementations of the Refine and Coarsen algorithms.
5. Statistical Learning Approaches
In this section, we introduce two new algorithms that address the minimal NAP specification problem. While they share the same overarching concept as refine and coarsen, these methods employ sampling and statistical learning principles to efficiently learn a minimal NAP specification. Recall that the set of mandatory neurons $M$ is the union of all neurons that appears in any minimal specification and is an upper bound of the largest minimal specification.
5.1. The Statistical Refine Approach
Mandatory neurons are crucial for forming NAP specifications, as their binary states play a critical role in determining the neural network’s robustness performance. We leverages this property to find mandatory neurons statistically. To be more specific, suppose we sample some NAPs $P_{1},P_{2},\ldots,P_{n}$ from the NAP family $\mathcal{P}$. For those NAPs that qualify as specifications (i.e., $\mathcal{V}(P)=1$), mandatory neurons should appear more frequently in them than in those NAPs that fail the verification tool (i.e., $\mathcal{V}(P)=0$).
Based on this insight, we propose an approach called Sample_Refine that relies on non-repetitive sampling to identify mandatory neurons for solving the minimal NAP problem. We start with the coarsest NAP and iteratively collect the most probable mandatory neurons. In every iteration, we sample $k$ NAPs by refining unvisited neurons with some probability $\theta$. The sampled NAPs are then fed to the verification tool, and the neuron that appears most frequently in verifiable NAPs is the most probable mandatory neuron in this iteration. The neuron is marked as visited, and the process stops when either we collect $s$ neurons (assuming that $s$ is known), or the current mandatory neurons form a NAP specification $P$, i.e., $\mathcal{V}(P)=1$. Finally, we return the learned NAP, obtained by applying $\widetilde{\mathcal{A}}$ to the collected neurons. The algorithm 5 provides an overview of the above procedure.
It is worth noting that Sample_Refine doesn’t guarantee correctness, as we may end up collecting only the $s$ most probable mandatory neurons, which may not be sufficient to form a specification. Another concern regarding this algorithm is sampling efficiency, specifically the potential for the number of samples required to grow exponentially with the size of the minimal NAP specification $s$. To understand why this is problematic, consider a scenario where the only minimal NAP specification $P$ consists of all mandatory neurons; in this case, all $|M|$ neurons must be selected for $P$ to be learned. If $\theta$ is set to a constant value, then the expected number of samples needed to obtain the NAP specification is $(\frac{1}{\theta})^{|M|}$. To address this, we set $\theta=\left(\frac{|M|}{|M|+1}\right)^{|M|}$. This choice ensures that the sampling efficiency is polynomial in both $|M|$ and $s$, as proven in Theorem 5.1. Please refer to the proof in the Appendix B. In addition, Theorem 5.1 also shows that with high probability, a mandatory neuron will be found with $O(log|N|)$ calls to $\mathcal{V}$.
Theorem 5.1.
With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
5.2. The Statistical Coarsen Approach
The Coarsen algorithm begins with the most refined NAP and progressively coarsens each neuron until the verification process fails. Enhancing the algorithm’s performance is possible by coarsening multiple neurons during each iteration. However, a fundamental question emerges: How do we determine which set of neurons to coarsen in each round?
We present Sample_Coarsen to answer this question. In this approach, we assume that each neuron is independent of the others and select neurons to coarsen in a statistical manner. Specifically, in each iteration, we randomly coarsen a subset of refined neurons in the current NAP simultaneously to see if the new NAP can pass verification. We repeat this process until the NAP size reaches $s$. Algorithm 6 provides the pseudocode for Sample_Coarsen.
Input: The neural network $N$, the probability $\theta$, sample size $k$, and the size $s$
Output: A minimal NAP specification $P$
1 Function Sample_NAPs($unvisited$, $\theta$)
       $P\leftarrow\dot{\mathcal{A}}(N)$   /* Use the coarsest NAP as a blank template */
2       for $N_{i,l}$ in $unvisited$ do
3             $rand\leftarrow\text{random}(0,1)$ if $rand\leq\theta$ then
                   $P\leftarrow\widetilde{\Delta}(N_{i,l})$   /* Refine unvisited neurons using $\widetilde{\Delta}$ with probability $\theta$/
4
5
6      return $P$
7Function Sample_Refine($visited$, $\theta$, $k$, $s$)
8       while $|visited|<s\text{ and }\mathcal{V}(P)==0$  do
             $unvisited\leftarrow N\setminus visited$ $ctr=dict()$   /* Create a counter for each neuron in $unvisited$/
9             for _normal-_\__ in $range(k)$ do
                   $P\leftarrow Sample\_NAPs(unvisited,\theta)$   /* Sample $k$ NAPs */
10                   for $N_{i,l}$ in $unvisited$ do
11                         if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ and $\mathcal{V}(P)==1$ then
                               $ctr[N_{i,l}]$ += 1   /* Count successful NAPs for each neuron */
12
13
14
            $N_{mad}\leftarrow\operatorname*{argmax}_{N_{i,l}\in unvisited}\{ctr[N_{i,l}]\}$   /* Pick potential mandatory neuron */
15             $visited\leftarrow visited\cup\{N_{mad}\}$
16      return $\widetilde{\mathcal{A}}(visited)$
[…]
       return $None$   /* Return None if the most refined NAP fails verification */
18
19 else
20       if heuristics then
21             $visited\leftarrow Gradient\_Search(N)\cap Adversarial\_Prune(N)$
22       else
             $visited\leftarrow\emptyset$   /* Start from the coarsest NAP */
23
24      $Sample\_Refine(visited,\theta,k,s)$
Algorithm 5 Sample_Refine
Similar to Sample_Refine, Sample_Coarsen also faces the same challenge related to sample efficiency. To better illustrate this, suppose a minimal NAP specification $P$ with size $s$ can be found after one iteration of refinement. Then, the probability of selecting the exact number of $s$ mandatory neurons in $P$ is $\theta^{s}$. Consequently, if $\theta$ is set to a constant value, the expected number of samples needed to find the NAP is $(\frac{1}{\theta})^{s}$. Unlike Sample_Refine, we know the refinement will always ensure the resulting NAP can pass verification. So, once such a NAP is learned, we can narrow down the estimated mandatory neurons by an expected factor of $\theta$. In this way, the expected number of samples and the expected number of iterations are inversely related, yet their product is the total number of calls to $\mathcal{V}$. We demonstrate that by setting $\theta=e^{-\frac{1}{s}}$, we not only make the expected number of samples polynomial in $s$, but also minimize the total number of calls to $\mathcal{V}$, as proven in Theorem 5.2. Please refer to the proof in Appendix B.
Theorem 5.2.
With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
Input: The neural network $N$, the probability $\theta$, and the size $s$
Output: A minimal NAP specification $P$
1 Function Sample_Coarsen($mand\_neurons$, $\theta$, $s$)
2       $P\leftarrow\widetilde{\mathcal{A}}(N)$; $mand\_neurons\leftarrow N$ if $\mathcal{V}(P)==0$ then
             return $None$   /* Return None if the most refined NAP fails verification */
3
4       else
5             while $|P|>s$ do
6                   $P\leftarrow Sample\_NAPs(mand\_neurons,\theta)$ if $\mathcal{V}(P)==1$  then
7                         $found\_neurons\leftarrow\emptyset$ for $N_{i,l}$ in $N$ do
8                               if $P_{i,l}==\widetilde{\mathcal{A}}(N_{i,l})$ then
9                                     $found\_neurons\leftarrow found\_neurons\cup\{N_{i,l}\}$
10
                        $mand\_neurons\leftarrow found\_neurons$   /* Reduce search space */
11
12                   else
                         $P\leftarrow Sample\_Naps(mand\_neurons,\theta)$   /* Sample a new NAP */
13
14
            return $P$   /* Return the minimal NAP of size $s$/
15
16
Algorithm 6 Sample_Coarsen
Setting the sample probability $\theta$
Setting $s$ poses a challenge in practice, as we assume that $s$ is always provided in Sample_Refine and Sample_Coarsen. However, this can be addressed by dynamically updating $\theta$ based on the result of $\mathcal{V}(P)$ (Liang et al., 2011). With $\theta$ from theorem 5.2, Sample_Coarsen finds a NAP specification with probability $\left(e^{-1/s}\right)^{s}=e^{-1}$. Recall that theorem 5.1 state that Sample_Refine finds a NAP specification with probability $\left(\frac{|M|}{|M|+1}\right)^{|M|}$. The probability approaches $e^{-1}$ asymptotically. Thus, we aim to set $\theta$ such that the $Pr(\mathcal{V}(P)=1)=e^{-1}$. Intuitively, if a sampled NAP $P$ is a specification, we decrease $\theta$ so less neurons will be refined or more neurons will be coarsened. Similarly, if $P$ is not a specification, $\theta$ needs to be increased.
Given that $\theta\in[0,1]$, we can parameterize it using the Sigmoid function $\sigma(\lambda)=\left(1+e^{-\lambda}\right)^{-1}$, where $\lambda\in(-\infty,\infty)$. Since $Pr(\mathcal{V}(P)=1)$ depends on $\theta$ as well, we express it as a function of $\lambda$, $g(\lambda)=Pr(\mathcal{V}(P)=1)$. Then, setting $Pr(\mathcal{V}(P)=1)=e^{-1}$ can be achieved through the following minimization problem:
(14) $L(\lambda)=\frac{1}{2}(g(\lambda)-e^{-1/s})^{2}$
The loss function $L(\lambda)$ can be minimized by statistical learning using stochastic gradient descent. With a step size $\eta$, update $\lambda$ using $\lambda\leftarrow\lambda-\eta\frac{dL}{d\lambda}$. Note that $\frac{dL}{d\lambda}$ can be expressed as:
(15) ${\frac{dL}{d\lambda}}=(g(\lambda)-e^{-1/s})\frac{dg(\lambda)}{d\lambda}$
Given $g(\lambda)=Pr(\mathcal{V}(P)=1)$, we can replace $g(\lambda)$ with $\mathcal{V}(P)$ for stochastic gradient update. Additionally, since $\frac{dg(\lambda)}{d\lambda}>0$, we simply ignore it as its multiplication effect can be represented by $\eta$. Therefore, the final update rule is given by:
(16) $\lambda\leftarrow\lambda-\eta(\mathcal{V}(P)-e^{-1})$
6. Volume Estimation of $R_{P}$
Figure 7. Volume Estimation of $R_{P}$ using an orthotope in a 2-dimensional case. The gray rectangle represents the input space, with the training set depicted by a collection of data points $+$. The polygon corresponds to some NAP $P$. Initially, we identify an anchor point, denoted by $+$. Then, we construct the orthotope, represented by the green rectangle, by extending upper and lower bounds starting from the anchor point until it extends beyond $P$.
Conceptually, NAP specifications typically correspond to significantly larger input regions compared to local neighbourhood specifications. This serves as the primary motivation for utilizing NAPs as specifications. However, previous work lacks sufficient justification or evidence to support this claim. In this section, we propose a simple method for approximating the volume of $R_{P}$, i.e., the region corresponding to a NAP $P$. This allows us to: 1) quantify the size difference between $R_{P}$ and $L_{\infty}$ ball specifications; 2) gain insights into the volumetric change from the most refined NAP specification to the minimal NAP specification.
Computing the exact volume of $R_{P}$ is at least NP-hard, as determining the exact volume of a polygon is known to be NP-hard (Dyer and Frieze, 1988). Moreover, computing the exact volume of $R_{P}$ can be even more challenging due to its potential concavity. To this end, our method estimates the volume of $R_{P}$ by efficient computation of an orthotope that closely aligns with $R_{P}$, as illustrated in Figure 7. We briefly describe it as follows:
Finding an anchor point
The first step is to find an anchor point to serve as the center of the orthotope. Ideally, this anchor point should be positioned close to the center of $R_{P}$ to ensure a significant overlap between the orthotope and $R_{P}$. However, computing the actual center of $R_{P}$ is costly. Thus, we look for a pseudo-center from the training set $X$ that resides in $R_{P}$. This pseudo-center can be computed by finding the point that uses the smallest $L_{\infty}$ ball to cover other data points, solved as the following optimization problem:
$c_{\text{pseudo}}=\underset{x\in R_{P}}{\arg\min}\max_{x^{\prime}\in R_{P}}\|x-x^{\prime}\|_{\infty}$
where $R_{P}=\{x\mid\mathcal{A}(N,x)\preccurlyeq P,x\in X\}$. When $|X|$ is small, $c_{\text{pseudo}}$ can be computed directly; for larger $|X|$, a statistical computation strategy is required.
Constructing the orthotope
Once the pseudo-center $c_{\text{pseudo}}$ is determined, we want to create an orthotope around $c_{\text{pseudo}}$ to closely align with $R_{P}$. The orthotope is constructed by determining pairs of upper and lower bounds $U^{(i)}$ and $L^{(i)}$ for each dimension $i$. Specifically, $U^{(i)}$ and $L^{(i)}$ are computed through expansion in two opposite directions from $c_{\text{pseudo}}$ along dimension $i$ until they extend beyond $R_{P}$. This expansion can be expressed as:
$\max_{U^{(i)}}\{x^{\prime}\in R_{P}\,|\,x^{\prime}:=c_{\text{pseudo}}+U^{(i)}\}\text{ ; }\max_{L^{(i)}}\{x^{\prime}\in R_{P}\,|\,x^{\prime}:=c_{\text{pseudo}}-L^{(i)}\}$
Here, $U^{(i)}$ and $L^{(i)}$ represent the upper and lower bounds in dimension $i$ respectively, originating from $c_{\text{pseudo}}$. These bounds can be efficiently calculated with binary search.
The choice of the archer point is crucial in our approach. If it is located at a corner of $R_{p}$, the volume calculation will be highly biased. This can pose a problem when we seek to understand the volumetric change from the most refined NAP specification to the minimal NAP specification. Additionally, using the orthotope as an estimator provides convenience in understanding the volumetric change simply by examining differences in each input dimension.
7. Evaluation
In this section, we conduct a comprehensive evaluation of our methods for learning minimal NAP specifications across a range of benchmarks, spanning from a simple binary classifier to the state-of-the-art image classification model. We illustrate that minimal NAP specifications usually involve only a fraction of neurons compared to the most refined NAPs, yet they dramatically extend the verifiable bounds to several magnitudes larger.
Experiment Setup
All experiments in this section were conducted on an Ubuntu 20.04 LTS machine with 172 GB of RAM and an Intel(R) Xeon(R) Silver Processor. For verification, we utilized Marabou (Katz et al., 2019), a dedicated state-of-the-art neural network verifier. We configured a timeout of 10 minutes for each call to the verification tool. If the timeout is exceeded, the current neuron is retained in the minimal NAP specification even if its status cannot be determined.
[…]
0 1
$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
The $\widetilde{\mathcal{A}}$ function 102 - 78.11 81.40 93 - 87.06 80.28
Coarsen 31 102 98.22 95.35 32 93 99.65 94.37
Adversarial_Prune 61 - 81.06 88.37 54 - 87.06 80.28
Gradient_Search 43 - 82.84 88.37 39 - 93.37 87.32
Sample_Refine 53 187 92.43 89.37 56 264 89.86 88.51
Sample_Coarsen 42 47 94.69 92.34 45 41 94.15 91.52
Wisconsin Breast Cancer with Binary Classifier
We conduct our first experiment using a four-layer neural network as a binary classifier, where each layer consists of 32 neurons. This classifier is trained on the Wisconsin Breast Cancer (WBC) dataset (William et al., 1995), representing a decision-critical task where robustness is essential. Our trained model achieves a test set accuracy of 95.61%. We calculate the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ for labels 0 and 1 using the statistical abstraction function $\widetilde{\mathcal{A}}$ with a confidence ratio of $\delta=0.95$. The size of $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ are 102 and 93, respectively. In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0 and 1 are significantly reduced to 31 and 32, respectively. It is worth mentioning that our estimation approaches provide a fairly accurate estimate of the mandatory neurons, despite computing rather loose upper bounds. To be more specific, Adversarial_Prune and Gradient_Search compute 61 and 43 mandatory neurons for label 0, respectively. Together, they cover 25 out of the 31 mandatory neurons appearing in the minimal NAP specification for label 0. For label 1, Adversarial_Prune and Gradient_Search compute 54 and 39 mandatory neurons, respectively, covering 25 out of the 32 mandatory neurons appearing in the minimal NAP specification for label 1.
Regarding the statistical approaches, Sample_Refine computes NAP specifications of size 53 and 56 for label 0 and label 1, respectively. This is achieved while making 187 and 264 calls, respectively. Although these numbers are quite large, it is expected since it has to sample multiple trials in each iteration. In contrast, Sample_Coarsen is significantly more efficient. It can learn NAP specifications of size 42 and 45 for label 0 and label 1 using only 47 and 41 calls, respectively.
Recall that one of the main motivations for learning the minimal NAP specifications is their potential to verify larger input regions compared to the refined NAP specifications. To support this claim, we compute the percentile of unseen test data that can be verified using these NAPs. We choose test data as they are sampled from the input space and can be viewed as a proxy to understand the verifiable bounds of different NAPs. We observe that the most refined NAP specifications $\widetilde{P}^{0}$ and $\widetilde{P}^{1}$ can cover 81.40% and 80.28% of the test data for labels 0 and 1, respectively. In contrast, this coverage significantly increases to 95.35% and 94.37% of the test data for the minimal NAP specifications of labels 0 and 1, respectively. To intuitively understand the extent of change in verifiable regions from the refined NAP specifications to the minimal NAP specifications, we also compare their estimated volume of the corresponding region $R_{P}$. The estimated volumetric changes are substantial, with label 0 experiencing an increase 58324.35 times larger and label 1 showing an increase 3053.53 times larger, respectively. To comprehend this change in each dimension, Figure 8 illustrates the comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1, referencing the anchor point.
(a) Comparing verifiable input ranges for label 0
(b) Comparing verifiable input ranges for label 1
Figure 8. Comparison of verifiable input ranges between the refined NAP specifications and the minimal NAP specifications for labels 0 and 1 on the Wisconsin Breast Cancer dataset, with reference to the anchor point. The minimal NAP specifications enable verification over a larger input range for all dimensions.
Table 2. Overview of the size of learned minimal NAP specifications on the MNIST benchmark.
0 1 4
$P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test $P$$\mathcal{V}$ train test
The $\widetilde{\mathcal{A}}$ function 751 - 79.50 80.51 745 - 86.01 85.11 712 - 77.54 80.24
Coarsen 480 751 98.68 98.78 491 745 98.90 98.59 506 712 98.51 97.45
Adversarial_Prune 618 - 83.13 71.32 630 - 79.02 76.51 699 - 82.66 84.12
Gradient_Search 195 - 92.1 91.80 237 - 89.73 88.39 169 - 87.59 88.51
Sample_Refine 564 4059 90.03 87.91 570 3663 88.36 86.84 581 4532 82.90 83.46
Sample_Coarsen 532 33 93.19 93.01 559 27 94.12 93.68 562 25 93.89 93.64
MNIST with Fully Connected Network
To show that our insights and approaches can be applied to more complicated datasets and networks, we conduct the second set of experiments using the mnistfc_256x4 model (VNNCOMP, 2021), a 4-layer fully connected network with 256 neurons per layer trained on the MNIST dataset. We specifically focus on classes 0, 1, and 4, as the most refined NAPs for other labels either encounter timeouts or fail the verification. We compute the most refined NAP specifications $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ for labels 0, 1, and 4 using the $\widetilde{\mathcal{A}}$ function with a confidence ratio of $\delta=0.99$. The size of $\widetilde{P}^{0}$, $\widetilde{P}^{1}$, and $\widetilde{P}^{4}$ are 751, 745 and 712, respectively, which aligns with the results from (Geng et al., 2023).
In contrast, the sizes of the minimal NAP specifications learned by the Coarsen algorithm for labels 0, 1, and 4 are significantly reduced to 480, 491, and 506, respectively. Moreover, our estimation approaches provide highly accurate estimates of the mandatory neurons present in minimal NAP specifications. For instance, considering label 0, Adversarial_Prune and Gradient_Search identify 618 and 195 mandatory neurons, respectively. And 445 and 160 out of the 480 mandatory neurons presented in the minimal NAP specification are discovered by Adversarial_Prune and Gradient_Search, respectively.
[…]
Moreover, the learned minimal NAP specifications correspond to significantly larger verifiable regions compared to the refined NAP specifications. Using the percentile of test data as a metric, the minimal NAP specifications increase the coverage ratio from 80.51% to 98.78%, 85.11% to 98.59%, and 80.24% to 97.45% for labels 0, 1, and 4, respectively. From the perspective of estimated volume, the volumetric changes are on the order of $10^{8}$ times larger.
ImageNet with Deep Convolutional Neural Networks
Current verification approaches struggle to scale to state-of-the-art neural networks such as deep convolutional neural networks, making them inapplicable for verifying certain critical systems like deep learning-based autonomous driving software. Consequently, our verification-dependent approaches also face scalability challenges. However, we can still assess the network’s robustness from a neuron-centric perspective using our estimation approaches, an aspect not well-studied in the current literature.
Thus, for the third experiment, we choose a deep convolutional network, specifically VGG-19 trained on the ImageNet dataset, as the benchmark for estimating mandatory neurons. We select neurons from fully connected layers [^1] We leave NAP computation for other layers such as convolutional layers for future work. of VGG-19, which contains a total of 8192 neurons, to compute NAPs. The original ImageNet includes 1000 classes with millions of images. We also narrow our focus to the top five largest classes in ImageNet. Each class consists of around 1000 training images and 300 test images. Table 3 presents the results of the learned NAP using verification-free approaches.
To show that these estimated mandatory neurons are indeed meaningful, we employ several approaches from the model interpretability domain such as Grad-CAM (Selvaraju et al., 2017) and saliency maps (Simonyan et al., 2014a). We observe a strong correlation between neurons exhibiting high interpretability under certain metrics and our estimated mandatory neurons computed by Adversarial_Prune and Gradient_Search. However, we also find that some mandatory neurons do not seem to participate significantly in interpretability, yet they are indeed necessary for the robustness of models. We find this interesting and plan to investigate those neurons in future work.
Although we cannot definitively determine if these NAPs can serve as specifications, these NAPs do cover significant portions of unseen test data. For example, the NAP formed by mandatory neurons learned through Gradient_Search can cover $87.01\%$ of unseen test data. This aspect makes them a promising basis for computing NAP specifications, especially if verification tools become accessible for such large models in the future.
Table 3. Overview of the size of learned minimal NAP specifications on the ImageNet benchmark.
box_turtle spiny_lobster acorn_squash confectionery stone_wall
$P$ test $P$ test $P$ test $P$ test $P$ test
The $\widetilde{\mathcal{A}}$ function 1978 39.42 1006 39.61 1003 29.23 878 33.84 971 31.54
Adversarial_Prune 1863 39.42 955 39.61 823 25.68 845 23.07 865 39.61
Gradient_Search 611 53.30 572 87.01 301 52.40 256 44.23 260 54.90
8. Related Work
8.1. Neural Network Verification
Neural network verification has attracted much attention due to the increasingly widespread applications of neural networks in safety-critical systems. it’s NP-hard nature resulting from the non-convexity introduced by activation functions (Katz et al., 2017) makes it a challenging task. Thus most of the existing work on neural network verification focuses on designing scalable verification algorithms. For instance, while initially proposed solver-based approaches (Huang et al., 2017b; Ehlers, 2017; Cheng et al., 2017; Tjeng et al., 2019) were limited to verifying small-scale neural networks with fewer than 100 neurons, state-of-the-art approaches (Xu et al., 2021; Wang et al., 2021c; Lu and Kumar, 2019) can verify more complex neural networks. It is worth mentioning that most existing work adopts local neighborhood specifications to verify the robustness properties of neural networks (Shapira et al., 2023). Despite being a reliable measure, using specifications that define local neighborhood of reference data points may not cover any test data, let alone generalizing to the verification of unseen test set data. Yet evidently, being able to verify the neural network on test data is an important task in real-world applications. The development of NAP specification (Geng et al., 2023) sheds light on this challenge. While previous work on NAP specifications proves its effectiveness, it assumes that all neurons contribute equally to the model’s robustness. In this paper, we challenge this assumption by demonstrating that only a small fraction of neurons are crucially related to robustness. This not only allow us study robustness from the neurons’ perspective but also significantly pushes the known verifiable boundaries.
8.2. Abstract Interpretation
Abstract interpretation (Cousot and Cousot, 1977) is a fundamental concept in software analysis and verification, particularly for approximating the semantics of discrete program states. By sacrificing precision, abstract interpretation typically enables scalable and faster proof finding during verification (Cousot and Cousot, 2014). Although abstract interpretation for neural network verification has been proposed and studied in previous literature (Gehr et al., 2018; Mirman et al., 2018), abstract interpretation of neural activation patterns for verification is a relatively new field. Perhaps the most related work from the perspective of abstract interpretation is learning minimal abstractions (Liang et al., 2011). While our work shares similarities in problem formulation and statistical approaches, we address fundamentally different problems. One limitation in our work is that our abstraction states may be too coarse: value in range $(0,+\infty)$ is abstracted into one state. This approach could over-approximate neuron behavior and thus fail to prove certain properties. We observe that neuron values exhibit different patterns in range for different input classes, suggesting the potential existence of more abstraction states. We plan to explore this direction in our future work.
8.3. Neural Activation Patterns
Neural activation patterns have commonly been used to understand the internal decision-making process of neural networks. One popular line of research is feature visualization (Yosinski et al., 2015; Bäuerle et al., 2022), which investigates which neurons are activated or deactivated given different inputs. This is also naturally related to the field of activation maximization (Bau et al., 2017; Simonyan et al., 2014b; Erhan et al., 2009), which studies what kind of inputs could mostly activate certain neurons in the neural network. In this way, certain prediction outcomes may be attributed to the behavior of specific neurons, thereby increasing the interpretability of the underlying models. In summary, most of existing works focus on learning statistical correlations between either NAPs and inputs or NAPs and prediction outcomes. However, whether such correlations can be trusted or even verified is not fully studied. To this end, we propose the concept of mandatory neurons and highlight their importance in the robustness of model predictions. Such causal links between neurons and prediction outcomes are not only identified but also verified. We believe this ”identify then verify” paradigm can be extended to existing research on NAPs to certify our understanding of neural networks. And we leave the exploration of this direction for our future work.
9. Conclusion
We introduce a new problem — learning the minimal NAP specification, and discuss its importance in neural network verification. Finding minimal NAP specification not only enables the verification of larger input regions compared to existing methods but also provides a means of inspecting when and how neural networks can make reliable and robust predictions. To solve this problem, we first propose two simple approaches Refine and Coarsen, which leverage off-the-shelf verification tools that find the minimal NAP specification with correctness guarantees. We also propose the statistical version of Refine and Coarsen, that combine sampling and statistical learning principles to achieve correctness in an more efficient manner. However, these approaches depend on underlying verification tools that are computationally expensive. To this end, we propose two approximate approaches, Adversarial_Prune and Gradient_Search. These methods utilize adversarial attacks and local gradient computation to efficiently investigate potential causal relationships between specific neurons and the model’s robustness. This is a task that current verification frameworks struggle to accomplish. Finally, to appreciate the volumetric change from the most refined NAP specification to the minimal NAP specification, we propose a simple method for estimating the volume of the region corresponding to a NAP. Our experiments indicate that minimal NAP specifications utilize much smaller fractions of neurons compared to the most refined NAPs. Nevertheless, they enable a substantial expansion of the verifiable boundaries by several orders of magnitude.
Appendix A Proofs of simple approaches
Theorem A.1 (Property of Refine).
The algorithm Refine returns a minimal NAP specification with $\mathcal{O}(2^{|N|})$ calls to $\mathcal{V}$.
Proof.
Let $P$ be the returned NAP. We prove this by contradiction. Suppose we can further refine $P$, meaning there exists a NAP $P^{\prime}$ such that $|P^{\prime}|\leq|P|$ and $\mathcal{V}(P^{\prime})=1$. However, the algorithm states that any $P^{\prime}$ smaller than $|P|$ fails verification, which contradicts $\mathcal{V}(P^{\prime})=1$.
In the worst case, the NAP size $k$ runs up to $|N|$. For each $k$, we need to check $\binom{|N|}{k}$ number of NAPs. In total, this number of NAPs we need to check is $\binom{|N|}{1}+\binom{|N|}{2}+\cdots+\binom{|N|}{|N|}=2^{|N|}-1$ according to the binomial theorem, resulting in a runtime complexity of $\mathcal{O}(2^{|N|})$.
Theorem A.2 (Property of Coarsen).
The algorithm Coarsen returns a minimal NAP specification with $\mathcal{O}(|N|)$ calls to $\mathcal{V}$.
Proof.
Let $P$ be the NAP returned by Coarsen. Our goal is to show that any $P^{\prime}$ smaller than $P$ results in $\mathcal{V}(P^{\prime})=0$. To construct such a smaller $P^{\prime}$, we need to apply the refine action $\widetilde{\Delta}$ on $P$ through some neuron $N_{i,l}$, i.e., $P^{\prime}:=\widetilde{\Delta}(N_{i,l})=\widetilde{\mathcal{A}}(N_{i,l})\cup P\setminus P_{i,l}$. According to the algorithm, $\mathcal{V}(P^{\prime})=0$. In the worst case, the algorithm needs to iterate through each neuron in $N$, resulting in a runtime complexity of $\mathcal{O}(|N|)$. ∎
Appendix B Proofs of statistical approaches
Our proofs of properties of Sample_Refine and Sample_Coarsen mainly follow those in (Liang et al., 2011). Interested readers may refer to it for detailed proofs.
Theorem B.1 (Property of Sample_Refine).
With probability $\theta=|(\frac{|M|}{|M|+1})|^{|M|}$, Sample_Refine has $1-\delta$ probability of outputting a minimal NAP specification with $\Theta\left(|M|^{2}(\log|N|+\log(s/\delta))\right)$ examples each iteration and $O(s|M|^{2}(\log|N|+\log(s/\delta)))$ total calls to $\mathcal{V}$.
Proof.
Considering the size of the minimal specification $s$, Sample_Refine will execute $s$ iterations. If we sample $k$ times in each iteration, then the probability of selecting a mandatory neuron is at least $1-\frac{\delta}{s}$. Consequently, by applying a union bound, the algorithm will identify a NAP specification with a probability of at least $1-\delta$.
Now, let’s delve deeper into one iteration of the process. The fundamental concept is that a mandatory neuron $m$ demonstrates a stronger correlation with proving the robustness query ($\mathcal{V}(P)=1$) compared to a non-mandatory one. This enhanced correlation increases the probability of its selection significantly when $k$ is sufficiently large.
Let’s revisit the notion of $M$, representing the set of mandatory neurons with a size of $|M|$. Now, let’s focus on a specific mandatory neuron $n^{+}\in M$. We define $B_{j-}$ as the event indicating $k_{n^{-}}>k_{n^{+}}$, and $B$ as the event where $B_{j-}$ holds for any non-mandatory neuron $n^{-}\in N\setminus M$. Importantly, if $B$ fails to occur, then the algorithm will correctly identify a mandatory neuron. Hence, our primary objective is to establish that $Pr(B)\leq\frac{\delta}{s}$. Initially, employing a union bound provides:
(17) $Pr(B)\leq\sum_{n^{-}}Pr(B_{n^{-}})\leq|N|\max_{n^{-}}Pr(B_{n^{-}})$
Let’s delve into each training example $P^{(i)}$ and introduce the notation $X_{i}=(1-\mathcal{V}(P^{(i)}))(P_{n^{-}}^{(i)}-P_{n^{+}}^{(i)})$. It’s worth emphasizing that $B_{n^{-}}$ manifests precisely when $\frac{1}{n}(k_{n^{-}}-k_{n^{+}})=\frac{1}{n}\sum_{i=1}^{n}X_{i}>0$. Our objective now is to bound this quantity utilizing Hoeffding’s inequality, considering the mean as:
(18) $\mathbb{E}[X_{i}]=Pr(\mathcal{V}(P)=1,P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})-Pr(\mathcal{V}(P)=1,P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$
and the bounds are $-1\leq\mathbb{E}[X_{i}]\leq 1$ Setting $\epsilon=-\mathbb{E}[X_{i}]$, we get:
(19) $Pr(B_{n^{-}})\leq e^{\frac{-k\epsilon^{2}}{2}},n^{+}\in M,n^{-}\notin M.$
Substituting (16) into (15) and rearranging terms, we can solve for $k$:
(20) $\frac{\delta}{s}\leq|N|e^{\frac{-k\epsilon^{2}}{2}}\text{ implies}k\geq\frac{2(log|N|+log(\frac{s}{\delta}))}{\epsilon^{2}}$
Our attention now shifts towards deriving a lower bound for $\epsilon$, which intuitively reflects the discrepancy (in terms of correlation with proving the robustness query) between a mandatory neuron and a non-mandatory one. It’s noteworthy that $Pr(P_{n}\in{\{\mathbf{1},\mathbf{0}}\})=\theta$ for any $n\in N$. Furthermore, since $n^{-}$ is non-mandatory, we can assert that $Pr(\mathcal{V}(P)=1|P_{n^{-}}\in\{\mathbf{1},\mathbf{0}\})=Pr(\mathcal{V}(P)=1)$. By leveraging these observations, we can express:
(21) $\epsilon=\theta(Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})-Pr(\mathcal{V}(P)=1))$
Let’s view $C$ as the collection of minimal NAP specifications. We can treat $C$ as a set of clauses in a Disjunctive Normal Form (DNF) formula: $\mathcal{V}(P;C)=\neg\bigvee_{c\in C}\bigwedge_{n\in c}P_{n}$, where we explicitly specify the dependence of $\mathcal{V}$ on the clauses $C$. For example, if $C={{1,2},{3}}$, it corresponds to $\mathcal{V}(P)=\neg[(P_{1}\land P_{2})\lor P_{3}]$. Now, let $C_{j}={c\in C:n\in c}$ denote the clauses containing $n$. We reformulate $Pr(\mathcal{V}(P)=1)$ as the sum of two components: one originating from the mandatory neuron $n^{+}$ and the other from the non-mandatory neuron:
(22) $Pr(\mathcal{V}(P)=1)$ $=Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
(23) $\quad+Pr(\mathcal{V}(P;C\setminus C_{n^{+}})=1)$
Calculating $Pr(\mathcal{V}(P)=1|P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\})$ follows a similar process. The only distinction arises from conditioning on $P_{n^{+}}\in\{\mathbf{1},\mathbf{0}\}$, which introduces an extra factor of $\frac{1}{\theta}$ in the first term because conditioning divides by $Pr(P_{n^{+}})=\theta$. The second term remains unchanged since no $c\notin C_{n^{+}}$ is mandatory in $P_{n^{+}}$. Substituting these two outcomes back into the equation yields:
(24) $\epsilon=(1-\theta)Pr(\mathcal{V}(P;C_{n^{+}})=1,\mathcal{V}(P;C\setminus C_{n^{+}})=0)$
Now, our objective is to establish a lower bound for (22) across all possible $\mathcal{V}$ (equivalently, $C$), where $n^{+}$ is permitted to be mandatory in $C$. Interestingly, the worst possible $C$ can be obtained by either having $|M|$ disjoint clauses ($C={{n}:n\in M}$) or a single clause ($C={M}$ if $s=|M|$). The intuition behind this is that when $C$ consists of $|M|$ clauses, there are numerous possibilities ($|M|-1$ of them) for some $c\notin C_{n^{+}}$ to be true, making it challenging to determine that $n^{+}$ is a mandatory neuron; in such cases, $\epsilon=(1-\theta)\theta(1-\theta)^{|M|-1}$. Conversely, if $C$ comprises a single clause, then letting this clause be true becomes exceedingly challenging; in this scenario, $\epsilon=(1-\theta)\theta^{|M|}$.
Let’s consider the scenario where $C$ comprises $|M|$ clauses. We aim to maximize $\epsilon$ concerning $\theta$ by setting the derivative $\frac{d\epsilon}{d\theta}=0$ and solving for $\theta$. This optimization yields $\theta=\frac{1}{|M|+1}$ as the optimal value. Substituting this value into the formula of $\epsilon$, we obtain $\epsilon=\frac{1}{|M|+1}(\frac{|M|}{|M|+1})^{|M|}$. Note that $(\frac{|M|}{|M|+1})^{|M|}$ can be lower bounded by $e^{-1}$, implying that $\epsilon^{-2}=\mathcal{O}(|M|^{2})$. Substituting this result to equation (18) will conclude the proof. ∎
Theorem B.2 (Property of Sample_Coarsen).
With probability $\theta$=$e^{-\frac{1}{s}}$, Sample_Coarsen learns a minimal NAP specification with $\mathcal{O}(slog|N|)$ calls to $\mathcal{V}$.
Proof.
Let’s first estimate the number of calls that Sample_Coarsen makes to $\mathcal{V}$. We denote the number of calls as $\mathcal{C}(P^{L})$, where $P^{L}$ is the most refined NAP. Then $\mathcal{C}(P^{L})$ can be computed recursively using the following rule:
(25) $\mathcal{C}(P^{L})=\begin{cases}|P|&\text{ if }|P^{L}|\leq s+1\\ 1+\mathbb{E}[(1-\mathcal{V}(P))\mathcal{C}(P)+\mathcal{V}(P)\mathcal{C}(P^{L})]&\text{ otherwise }\\ \end{cases}$
where $P$ is the sampled NAP. By assumption, there exists a NAP $P^{S}\preccurlyeq P$ of size $s$ that passes the verification. Define $G(P)=\neg(P^{S}\preccurlyeq P)$, which is 0 when $P^{S}$ subsumes $P$, i.e., all mandatory neurons in $P^{S}$ shows up in the sampled $P$. This follows that $Pr(G(P)=0)=Pr(P^{S}\preccurlyeq P)=\theta^{s}$. Note that $G(P)\geq\mathcal{V}(P)$, as the NAP $P^{S}$ suffices to prove the robustness query. We can estimate the upper bound of $\mathcal{C}(P^{L})$ by replacing $\mathcal{V}$ with $G$:
(26) $\mathcal{C}(P^{L})$ $\leq 1+\mathbb{E}[(1-G(P))\mathcal{C}(P)+G(P)\mathcal{C}(P^{L})]$
(27) $\leq 1+\theta^{s}\mathbb{E}[\mathcal{C}(P)|P^{S}\preccurlyeq P]+(1-\theta^{s})\mathcal{C}(P^{L})$
(28) $\leq\mathbb{E}[\mathcal{C}(P)|P^{S}\preccurlyeq P]+\theta^{-s}$
We now denote $\mathcal{C}(n)=max_{|P|=n}\mathcal{C}(P)$ as the maximum over NAP of size $n$. Note that given $P^{S}\preccurlyeq P$, $|P|=s+N$ where $N$ is a binomial random variable with $\mathbb{E}(N)=\theta(n-s)$.
Using the bound $\mathcal{C}(n)\leq(1-\theta^{n})\mathcal{C}(n-1)+\theta^{n}\mathcal{C}(n)\theta^{-s}$, we can observe that $\mathcal{C}(n)\leq\frac{\theta^{-s}}{1-\theta^{n}}\cdot n$. In addition, when $n$ is large enough, $\mathcal{C}(n)$ is concave, then by use Jensen’s inequality:
(29) $\mathcal{C}(n)\leq\mathcal{C}(\mathbb{E}[s+N])+\theta^{-s}=\mathcal{C}(s+\theta(n-s))+\theta^{-s}$
To solve the recurrence, this gives us:
(30) $\mathcal{C}(n)\leq\frac{\theta^{-s}logn}{log\theta^{-1}}+s+1$
The equation above illustrates a tradeoff between reducing the number of iterations (by increasing $\log\theta^{-1}$) and reducing the number of samples (by decreasing $\theta^{-s}$). To minimize $\mathcal{C}(n)$, we need to set $\theta$ so that the gradient of $\mathcal{C}(n)$ w.r.t $\theta^{-1}$ is 00. This gives us: $\frac{sx^{s-1}}{\log x}-\frac{x^{s-1}}{\log^{2}x}=0$, solving this gives us $\theta=e^{-\frac{1}{s}}$. Consequently, the upper bound becomes $\mathcal{C}(n)=es\log n+s+1=O(s\log n)$. ∎